OpenStudy (anonymous):

Write a possible explicit rule for the nth term of the sequence: 5, 10, 20, 40, 80, ...

OpenStudy (amistre64):

a possible eh

OpenStudy (amistre64):

5, 10, 20, 40, 80, 5 10 20 40 an = 2an-1

OpenStudy (anonymous):

Thank you(:

OpenStudy (amistre64):

i forget what explicits spose to look like tho; but thats at least the pattern :)

OpenStudy (amistre64):

an = 2 an-1 an-1 = 2 an-2 an = 2.2an-2 an = 2.2.2an-3 an = 2^3 an-3 an = 2^r an-r when r=n we get an-n = a0 an = 2^n an-n an = 2^n a0

OpenStudy (amistre64):

a0 = 5 sooo an = 5*2^n

OpenStudy (anonymous):

awesome thanks